OpenStudy (anonymous):

how do you do this equation? -13/24=5/-6+c

OpenStudy (anonymous):

i need detailed answer like the work too

OpenStudy (anonymous):

(-13/24)=(-5/6)+c you add (-5/6) to both sides (-13/24)+(5/6)=c (7/24)=c

OpenStudy (anonymous):

thanks man!!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

wait oops it was -5/16!!!!!!!

OpenStudy (anonymous):

hmm hang on

OpenStudy (anonymous):

so it would just be (-13/24)+(5/16) So -11/48=c

OpenStudy (anonymous):

yay i actually go tthat right the first time !!

OpenStudy (anonymous):

Cool